# Fair Summarization: Bridging Quality and Diversity in Extractive Summaries

### v3

@@ Abstract
Fairness in multi-document summarization of user-generated content remains a critical challenge in natural language processing (NLP). Existing summarization methods often fail to ensure equitable representation across different social groups, leading to biased outputs. In this paper, we introduce two novel methods for fair extractive summarization: FairExtract, a clustering-based approach, and FairGPT, which leverages GPT-3.5-turbo with fairness constraints. We evaluate these methods using Divsumm summarization dataset of White-aligned, Hispanic, and African-American dialect tweets and compare them against relevant baselines. The results obtained using a comprehensive set of summarization quality metrics such as SUPERT, BLANC, SummaQA, BARTScore, and UniEval, as well as a fairness metric $F$, demonstrate that FairExtract and FairGPT achieve superior fairness while maintaining competitive summarization quality. Additionally, we introduce composite metrics (e.g., SUPERT+$F$, BLANC+$F$) that integrate quality and fairness into a single evaluation framework, offering a more nuanced understanding of the trade-offs between these objectives. This work highlights the importance of fairness in summarization and sets a benchmark for future research in fairness-aware NLP models.
@@ 1 Introduction
Early research focused on summarizing formal text sources such as news articles. However, with the rise of social media, attention has shifted to summarizing user-generated content, which is diverse in style and language. Social media platforms bring together users from varied backgrounds, introducing linguistic diversity through informal language, slang, and emojis. This diversity raises the challenge of ensuring fairness in summarization ensuring balanced representation of various social groups. In social media, where public opinion is shaped, fair summaries are essential to include different perspectives and avoid underrepresentation of one or more social groups.
In the context of social media, where millions of users contribute diverse perspectives, ensuring representation of this diversity in summaries becomes crucial. Social media platforms encompass a wide range of voices, including those from historically underrepresented or marginalized groups, making it essential that summarization methods capture this diversity fairly. Without proper representation, certain voices might be excluded or misrepresented, leading to biased summaries that skew public discourse (Binns,, 2017; Hutchinson and Mitchell,, 2018). The need for fairness in summarization is further heightened by the fact that user-generated content is often informal and marked by dialectal variations, requiring models to go beyond traditional summarization approaches (Pitsilis et al.,, 2018). Therefore, ensuring that all groups—across race, gender, and linguistic diversity—are fairly represented is critical for generating balanced summaries that reflect the diversity of public opinion (Dash et al.,, 2018).
Despite advancements, bias remains a concern in automated summarization (Dash et al.,, 2019; Jung et al.,, 2019; Keswani and Celis,, 2021; Olabisi et al.,, 2022) as most existing summarization methods focus on quality but fall short in optimizing fairness. This gap leads to the key question: if a summarization method is optimized for fairness, how does it affect the overall summary quality?
Previous studies suggest a trade-off between fairness and quality (Jung et al.,, 2019). Improving fairness can sometimes lower quality. While existing algorithms have made strides in balancing the two, none achieve perfect fairness.
ChatGPT-EXT (Zhang et al.,, 2023) FairGPT (Ours)
Table 1: Comparison of summaries by ChatGPT-EXT and FairGPT. Tweets from different groups are highlighted: Group 1 (e.g., White-aligned) and Group 2 (e.g., African-American).
@@
To illustrate the performance of fairness-aware summarization models, we compare summaries generated by ChatGPT-EXT (Zhang et al.,, 2023) and our proposed FairGPT model on a sample instance from Divsumm dataset (Olabisi et al.,, 2022). As shown in Table 1, FairGPT ensures equal representation of tweets from different groups, while ChatGPT-EXT shows a slight imbalance.
@@
  • We propose FairExtract, a fair clustering-based extractive summarization method that achieves perfect fairness and is evaluated against baseline models using standard and composite quality-fairness metrics.
  • We develop FairGPT, a large language model-based extractive summarization method that enforces fairness through equal representation and accurate content extraction using the longest common subsequence.
@@ 2 Related Work
The field of natural language processing (NLP) has increasingly focused on addressing bias and fairness, driven by the demand for equity in AI systems. Research has explored two key dimensions: intrinsic bias, stemming from text representations, and extrinsic bias, reflecting performance disparities across demographic groups (Han et al.,, 2023).
Early work on fairness in summarization (Shandilya et al.,, 2018; Dash et al.,, 2019) revealed that summaries often fail to represent source data fairly, even when source texts from different groups have similar quality. This led to the development of fairness-aware algorithms across various stages of summarization, including pre-processing, in-processing, and post-processing techniques. For example, Keswani and Celis, (2021) proposed a post-processing method to mitigate dialect-based biases. Olabisi et al., (2022) introduced the DivSumm dataset, focusing on dialect diversity in summarization and evaluating algorithms on fairness.
Recent work has explored bias related to the position of input data. Olabisi and Agrawal, (2024) studied position bias in multi-document summarization, showing that the order of input texts affects fairness. Similarly, Huang et al., (2023) analyzed clustering-based summarization models, which may introduce political or opinion bias, emphasizing the need for fair representation.
Fair clustering, another key technique, has also seen significant research. Chierichetti et al., (2017) introduced the concept of fairlets—small, balanced clusters that ensure fair representation across protected groups. Building on this, Chen et al., (2019) proposed proportional centroid clustering to eliminate biases in cluster-based models.
Further advancements include scalable techniques for fair clustering, such as the fair $k$-median clustering method (Backurs et al.,, 2019), and approaches that generalize fairness constraints across multiple protected groups (Bera et al.,, 2019). Esmaeili et al., (2020) extended this work to probabilistic fair clustering, offering solutions for uncertain group memberships.
In the domain of clustering methodologies, Micha and Shah, (2020) explored fairness in centroid clustering, while Li et al., (2020) proposed Deep Fair Clustering (DFC), which leverages deep learning to filter sensitive attributes, improving both fairness and performance. This underscores the growing importance of combining fairness with robust clustering methods in NLP tasks.
@@ 3 Task Formulation
In this context, each document is a tweet from a specific dialect group, which serves as an indicator of its social group. Traditionally, various metrics like ROUGE (Lin,, 2004) and BERTScore (Zhang et al.,, 2019) have been used to evaluate summary quality. However, our primary focus is on balancing both quality and fairness, particularly in terms of representing different social groups equitably. To measure fairness, we use the Representation Gap (RG) metric, as proposed by Olabisi et al., (2022). This metric captures how well the summary reflects the proportions of the original groups. A lower RG score indicates better balance and thus a fairer summary.
@@
$\text{RG}(\mathcal{S})=\frac{\max\{N_{1}(\mathcal{S}),N_{2}(\mathcal{S})\}-\min\{N_{1}(\mathcal{S}),N_{2}(\mathcal{S})\}}{k}.$ (1)
@@
Our analysis faces two key challenges: (1) While quality metrics improve with larger values, fairness improves with smaller Representation Gap (RG) values. (2) Quality and fairness metrics differ greatly in scale, making direct comparison difficult.
@@
Finally, we introduce composite metrics, such as SUPERT+F, BLANC+F, SummaQA+F, BARTScore+F, and UniEval+F, which are the averages of the normalized quality metrics (e.g., SUPERT (Gao et al.,, 2020), BLANC (Vasilyev et al.,, 2020), SummaQA (Scialom et al.,, 2019), BARTScore (Yuan et al.,, 2021), and UniEval (Zhong et al.,, 2022)) and the fairness score $F$, providing a balanced assessment of both quality and fairness.
@@ 4 Fair Extractive Summarizers
In this work, we introduce two novel methods for fair extractive summarization: FairExtract and FairGPT. FairExtract utilizes clustering techniques with fairlet decomposition to ensure diversity in summaries while maintaining high-quality representation across different groups. FairGPT, on the other hand, leverages large language models (LLMs) such as GPT-3.5, incorporating fairness constraints and the longest common subsequence (LCS) method to match and fairly select content from different groups. Both methods prioritize fairness and ensure equitable representation in the generated summaries. The implementation code for both methods is available.  [^1] https://github.com/PortNLP/FairEXTSummarizer
@@ 4.1 FairExtract: A Clustering-based Fair Extractive Summarization Method
    Embedding Documents: We begin by embedding each document (tweet) into a high-dimensional space (e.g., using a pretrained model such as BERT (Devlin et al.,, 2019)), capturing its semantic content in Euclidean space. This embedding enables us to compute meaningful distances between documents, which is crucial for clustering.
@@
    Fairlet Decomposition: To ensure fairness in the summarization process, we decompose the dataset into fairlets. A fairlet is the smallest set of documents that maintains proportional balance between two groups, $G_{1}$ and $G_{2}$ (Backurs et al.,, 2019). If the proportions of $G_{1}$ and $G_{2}$ are $g_{1}$ and $g_{2}$, respectively, where $\text{gcd}(g_{1},g_{2})=1$, a fairlet must contain exactly $g_{1}$ documents from $G_{1}$ and $g_{2}$ documents from $G_{2}$. This ensures that the composition of the fairlet reflects the required ratio between the two groups, maintaining fairness at the smallest possible scale. The decomposition aims to minimize the sum of Euclidean distances between documents within the same fairlet.
@@
For a formal representation of the process, see Algorithm 2 in Appendix A.1.
@@ 4.2 FairGPT: An LLM-based Fair Extractive Summarization Method
FairGPT leverages GPT-3.5-turbo to generate fair extractive summaries by selecting an equal number of sentences from different social groups. It applies fairness checks and uses the longest common subsequence (LCS) to match generated summaries with the original tweets. Below are the detailed steps:
  1. 1.
  2. 2.
  3. 3.
    Matching using Longest Common Subsequence (LCS): As GPT sometimes generates partial sentences, we apply LCS to match the generated summary with the closest original tweets. The full tweets corresponding to the longest common subsequences are added to the final summary.
  4. 4.
  5. 5.
For a formal representation of the process, see Algorithm 1.
@@
system:   "You are an extractive fair summarizer that follows the
          output pattern. A fair summarizer should select the same
          number of sentences from each group of people."
user:     "Please extract sentences as the summary.
          The summary should contain {L} sentences which means
          select {L/2} number of sentences from each group of people
          to represent the idea of all groups in a fair manner.
@@ 5.1 Dataset
The dataset used in this study is DivSumm (Olabisi et al.,, 2022), consisting of tweets from three ethnic groups—White-aligned, Hispanic, and African-American—across 25 topics, with 30 tweets per group per topic, totaling 2,250 tweets. This diversity allows us to evaluate the fairness of our summarization methods across varied social and cultural contexts.
Our model works with two groups at a time, so we explore three pairings: White-Hispanic, Hispanic-African American, and White-African American. Each pairing maintains proportional representation from both groups to ensure an equitable balance in the summarization process.
For our experiments, we formed 60 tweets per group pair (30 from each group) and generated a 6-tweet summary per pair, covering all 25 topics. This yielded 75 distinct summaries per model, allowing us to evaluate both fairness and quality comprehensively. A sample of the dataset is available in Appendix A.2.
@@ 5.2 Baseline Methods
TextRank: TextRank is an unsupervised graph-based ranking method used for extractive summarization (Mihalcea and Tarau,, 2004). This standard vanilla baseline approach uses a single aggregated set of randomized documents from all groups as input for summarization, without any pre-processing.
BERT-Ext: BERT-Ext is an extractive summarization model that utilizes pre-trained embeddings from BERT and k-means clustering to select sentences closest to the centroid as summaries (Miller,, 2019). Similar to the TextRank baseline, we implemented BERT-Ext vanilla method.
Cluster-Heuristic (Cluster-H): This method first partitions the input documents into group-based subsets before generating separate group summaries of length . These group-level summaries are shuffled, combined and then used to generate a final, unified summary (Dash et al.,, 2019; Olabisi et al.,, 2022). As summarization models, we use TextRank and BERT-Ext.
Cluster-Automatic (Cluster-A): In this attribute-agnostic approach, documents are clustered automatically into $m$ subsets, and corresponding summaries of length are generated. The summaries are concatenated and used to generate a final summary (Olabisi et al.,, 2022). As summarization models, we experiment with TextRank and BERT-Ext.
ChatGPT-EXT: This approach uses GPT-3.5 for extractive summarization by employing in-context learning and chain-of-thought reasoning to identify key sentences. It focuses on extracting salient content from documents to generate coherent summaries while maintaining the structure of the original text (Zhang et al.,, 2023).

### v5

@@ Abstract
Fairness in multi-document summarization of user-generated content remains a critical challenge in natural language processing (NLP). Existing summarization methods often fail to ensure equitable representation across different social groups, leading to biased outputs. In this paper, we introduce two novel methods for fair extractive summarization: FairExtract, a clustering-based approach, and FairGPT, which leverages GPT-3.5-turbo with fairness constraints. We evaluate these methods using Divsumm summarization dataset of White-aligned, Hispanic, and African-American dialect tweets and compare them against relevant baselines. The results obtained using a comprehensive set of summarization quality metrics such as SUPERT, BLANC, SummaQA, BARTScore, and UniEval, as well as a fairness metric $F$, demonstrate that FairExtract and FairGPT achieve superior fairness while maintaining competitive summarization quality. Additionally, we introduce composite metrics (e.g., SUPERT+$F$, BLANC+$F$) that integrate quality and fairness into a single evaluation framework, offering a more nuanced understanding of the trade-offs between these objectives. Our code is available online. [^1] https://github.com/PortNLP/FairEXTSummarizer
Fair Summarization: Bridging Quality and Diversity in Extractive Summaries
Sina Bagheri Nezhad, Sayan Bandyapadhyay, Ameeta Agrawal Department of Computer Science Portland State University, USA {sina5,sayanb,ameeta@pdx.edu}
@@ 1 Introduction
ChatGPT-EXT (Zhang et al., 2023) FairGPT (Ours)
Table 1: Comparison of summaries generated by ChatGPT-EXT and FairGPT. Tweets from different groups are highlighted: Group 1 (e.g., White-aligned) and Group 2 (e.g., African-American).
@@
Early research focused on summarizing formal text sources such as news articles. However, with the rise of social media, attention has shifted to summarizing user-generated content, which is diverse in style and language (Dash et al., 2019; Jung et al., 2019; Keswani and Celis, 2021; Olabisi et al., 2022). Social media platforms bring together users from varied backgrounds, introducing linguistic diversity through informal language, slang, and emojis. This diversity raises the challenge of ensuring fairness in summarization for a balanced representation of various social groups. In social media, where public opinion is shaped, fair summaries are essential to include different perspectives and avoid underrepresentation of one or more social groups as without proper representation, certain voices might be excluded or misrepresented. Therefore, ensuring that all groups—across race, gender, and linguistic diversity—are fairly represented is critical for generating balanced summaries that reflect the diversity of public opinion (Dash et al., 2018). In particular, the dialectal variations among White-aligned, Hispanic, and African-American groups not only reflect different linguistic styles but also embody distinct cultural expressions that influence how users communicate.
Despite advancements, bias remains a concern in automated summarization (Dash et al., 2019; Jung et al., 2019; Keswani and Celis, 2021; Olabisi et al., 2022) as most existing summarization methods focus on quality but fall short in optimizing fairness. Improving fairness can sometimes lower quality (Jung et al., 2019). This gap leads to the key question: if a summarization method is optimized for fairness, how does it affect the overall summary quality?
@@
To illustrate the performance of fairness-aware summarization models, we compare summaries generated by ChatGPT-EXT (Zhang et al., 2023) and our proposed FairGPT model on a sample instance from Divsumm dataset (Olabisi et al., 2022). As shown in Table 1, FairGPT ensures equal representation of tweets from different groups, while ChatGPT-EXT shows a slight imbalance.
@@
  • We propose FairExtract, a fair clustering-based extractive summarization method that achieves perfect fairness while preserving competitive summarization quality, as demonstrated through evaluations against baseline models using standard and composite quality-fairness metrics.
  • We develop FairGPT, a large language model-based extractive summarization method that enforces fairness through equal representation and accurate content extraction using the longest common subsequence, producing fair summaries without sacrificing competitive summarization quality.
@@ 2 Related Work
The field of NLP has increasingly focused on addressing bias and fairness, with research focused along two key dimensions: intrinsic bias, stemming from text representations, and extrinsic bias, reflecting performance disparities across demographic groups (Han et al., 2023).
Early work on fairness in summarization (Shandilya et al., 2018; Dash et al., 2019) revealed that summaries often fail to represent source data fairly, even when source texts from different groups have similar quality. This led to the development of fairness-aware algorithms across various stages of summarization, including pre-processing, in-processing, and post-processing techniques. For example, Keswani and Celis (2021) proposed a post-processing method to mitigate dialect-based biases. Olabisi et al. (2022) introduced the DivSumm dataset, focusing on dialect diversity in summarization and evaluating algorithms on fairness.
Recent work has explored bias related to the position of input data. Olabisi and Agrawal (2024) studied position bias in multi-document summarization, showing that the order of input texts affects fairness. Similarly, Huang et al. (2023) analyzed clustering-based summarization models, which may introduce political or opinion bias, emphasizing the need for fair representation.
Recent work highlights that large language models often reflect dominant Western cultural norms, resulting in cultural bias Tao et al. (2024). Liu et al. (2024) provided a taxonomy for culturally aware NLP that emphasizes the role of values, norms, and linguistic diversity. Moreover, Hershcovich et al. (2022) discussed cross-cultural challenges in NLP and advocate for strategies that integrate cultural insights into model development.
Fair clustering, another key technique, has also seen significant research. Chierichetti et al. (2017) introduced the concept of fairlets—small, balanced clusters that ensure fair representation across protected groups. Building on this, Chen et al. (2019) proposed proportional centroid clustering to eliminate biases in cluster-based models.
Further advancements include scalable techniques for fair clustering, such as the fair $k$-median clustering method (Backurs et al., 2019), and approaches that generalize fairness constraints across multiple protected groups (Bera et al., 2019). Esmaeili et al. (2020) extended this work to probabilistic fair clustering, offering solutions for uncertain group memberships.
In the domain of clustering methodologies, Micha and Shah (2020) explored fairness in centroid clustering, while Li et al. (2020) proposed Deep Fair Clustering (DFC), which leverages deep learning to filter sensitive attributes, improving both fairness and performance. This underscores the growing importance of combining fairness with robust clustering methods in NLP tasks.
@@ 3 Task Formulation
In this context, each document is a tweet from a specific dialect group, which serves as an indicator of its social group. Traditionally, various metrics like ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2019) have been used to evaluate summary quality. However, our primary focus is on balancing both quality and fairness, particularly in terms of representing different social groups equitably. To measure fairness, we use the Representation Gap (RG) metric, as proposed by Olabisi et al. (2022). This metric captures how well the summary reflects the proportions of the original groups. A lower RG score indicates better balance and thus a fairer summary.
@@
$\text{RG}(\mathcal{S})=\textstyle\frac{\max\{N_{1}(\mathcal{S}),N_{2}(\mathcal{S})\}-\min\{N_{1}(\mathcal{S}),N_{2}(\mathcal{S})\}}{k}.$ (1)
@@
At this point, we recognize two key challenges: (1) While quality metrics improve with larger values, fairness improves with smaller Representation Gap (RG) values. (2) Quality and fairness metrics differ greatly in scale, making direct comparison difficult.
@@
Finally, we introduce composite metrics, such as SUPERT+F, BLANC+F, SummaQA+F, BARTScore+F, and UniEval+F, which are the averages of the normalized quality metrics (e.g., SUPERT (Gao et al., 2020), BLANC (Vasilyev et al., 2020), SummaQA (Scialom et al., 2019), BARTScore (Yuan et al., 2021), and UniEval (Zhong et al., 2022)) and the fairness score $F$, providing a balanced assessment of both quality and fairness.
@@ 4 Fair Extractive Summarizers
In this work, we introduce two novel methods for fair extractive summarization: FairExtract and FairGPT. FairExtract utilizes clustering techniques with fairlet decomposition to ensure diversity in summaries while maintaining high-quality representation across different groups. FairGPT, on the other hand, leverages large language models (LLMs) such as GPT-3.5, incorporating fairness constraints and the longest common subsequence (LCS) method to match and fairly select content from different groups. Both methods prioritize fairness and ensure equitable representation in the generated summaries.
@@ 4.1 FairExtract: A Clustering-based Fair Extractive Summarization Method
    Embedding Documents: We begin by embedding each document (tweet) into a high-dimensional space (e.g., using a pretrained model such as BERT (Devlin et al., 2019)), capturing its semantic content in Euclidean space. This embedding enables us to compute meaningful distances between documents, which is crucial for clustering.
@@
    Fairlet Decomposition: To ensure fairness in the summarization process, we decompose the dataset into fairlets. A fairlet is the smallest set of documents that maintains proportional balance between two groups, $G_{1}$ and $G_{2}$ (Backurs et al., 2019). Assume the desired ratio of documents from $G_{1}$ to $G_{2}$ is $g_{1}:g_{2}$, where $g_{1}$ and $g_{2}$ are coprime (i.e., $\gcd(g_{1},g_{2})=1$). Then, a fairlet is defined as the smallest group of documents that exactly preserves this ratio, containing exactly $g_{1}$ documents from $G_{1}$ and $g_{2}$ documents from $G_{2}$. This ensures that the composition of the fairlet reflects the required ratio between the two groups, maintaining fairness at the smallest possible scale. The decomposition aims to minimize the sum of Euclidean distances between documents within the same fairlet.
@@
For a formal representation of the process, see Appendix A.1.
@@ 4.2 FairGPT: An LLM-based Fair Extractive Summarization Method
{adjustwidth}
0in-1in
@@
system:  "You are an extractive fair summarizer that follows the output pattern. A fair summarizer
          should select the same number of sentences from each group of people."
user:    "Please extract sentences as the summary. The summary should contain {L} sentences
          which means select {L/2} number of sentences from each group of people to represent
          the idea of all groups in a fair manner.
@@
FairGPT leverages an LLM generate fair extractive summaries by selecting an equal number of sentences from different social groups. It applies fairness checks and uses the longest common subsequence (LCS) to match generated summaries with the original tweets. Below are the detailed steps:
  1. 1.
  2. 2.
  3. 3.
    Matching using Longest Common Subsequence (LCS): As GPT sometimes extracts partial sentences, we apply LCS to match the generated summary with the closest original tweets. The full tweets corresponding to the longest common subsequences are added to the final summary.
  4. 4.
  5. 5.
For a formal representation of the process, see Algorithm 1.
@@ 5 Experimental Setup
Group Tweet
White-aligned Turns out not all White Castles are the same. Why do you push me away Chicago?!
African American "I mean I’m from Chicago. I’ll cheer for the Bears, but I’m a bigger 49ers fan."
White-aligned Nothing makes me happier than seeing the Bulls win _____ #ChicagoBasketball #Bullieve
White-aligned If you see on the news something about the Chicago Kitchen Clown Bandits, then it will be referring to me, my friend Eten, and I.
African American Truuu we tryna find sum to do too.. I dnt wanna b n Chicago if ain’t nobody here.
White-aligned Oh yeah.. I’m good. Hangin’ up here in Chicago today. :)
Hispanic You girls have a safe flight.! See you in Chicago (:
… (Dataset continues with more examples)
@@ 5.1 Dataset
The dataset used in this study is DivSumm (Olabisi et al., 2022), consisting of tweets from three dialect groups—White-aligned, Hispanic, and African-American—across 25 topics, with 30 tweets per group per topic, totaling 2,250 tweets.
Our model works with two groups at a time, so we explore three pairings: White-Hispanic, Hispanic-African American, and White-African American. Each pairing maintains proportional representation from both groups to ensure an equitable balance in the summarization process. Table 2 presents a sample of the dataset used in this study, containing tweets from different social groups about Chicago.
For our experiments, we formed 60 tweets per group pair (30 from each group) and generated a 6-tweet summary per pair, covering all 25 topics. This yielded 75 distinct summaries per model, allowing us to evaluate both fairness and quality comprehensively.
@@ 5.2 Baseline Methods
TextRank: TextRank is an unsupervised graph-based ranking method used for extractive summarization (Mihalcea and Tarau, 2004). This standard vanilla baseline approach uses a single aggregated set of randomized documents from all groups as input for summarization, without any pre-processing.
BERT-Ext: BERT-Ext is an extractive summarization model that utilizes pre-trained embeddings from BERT and k-means clustering to select sentences closest to the centroid as summaries (Miller, 2019). Similar to the TextRank baseline, we implemented BERT-Ext vanilla method.
Cluster-Heuristic (Cluster-H): This method first partitions the input documents into group-based subsets before generating separate group summaries of length . These group-level summaries are shuffled, combined and then used to generate a final, unified summary (Dash et al., 2019; Olabisi et al., 2022). As summarization models, we use TextRank and BERT-Ext.
Cluster-Automatic (Cluster-A): In this attribute-agnostic approach, documents are clustered automatically into $m$ subsets, and corresponding summaries of length are generated. The summaries are concatenated and used to generate a final summary (Olabisi et al., 2022). As summarization models, we experiment with TextRank and BERT-Ext.
ChatGPT-EXT: This approach uses GPT-3.5 for extractive summarization by employing in-context learning and chain-of-thought reasoning to identify key sentences. It focuses on extracting salient content from documents to generate coherent summaries while maintaining the structure of the original text (Zhang et al., 2023).